OpenStudy (anonymous):

How many ways is it possible to arrange five people in 5 seats? What about six people in six seats?

OpenStudy (anonymous):

5p5 = 5!/(5-5)! = 5!/0! 0! = 1 so 5! = 5*4*3*2*1 = 120 ways

OpenStudy (anonymous):

How many ways to put 1 person in 6 seats? After you done that, you got 5 persons left and 5 seats......